OpenStudy (anonymous):

V varies inversely with T and V = 12 when T = 6. Which equation shows this relationship? A. V = 72T B. v=72/t C. V = 2T D. v=2/t

OpenStudy (mathstudent55):

The general equation for inverse variation is: If y varies inversely with x, then \(y = \dfrac{k}{x} \)

OpenStudy (mathstudent55):

In your case, V varies inversely with T, so use V instead of y and T instead of x in the equation above. Keep k as it is. What do you get?

OpenStudy (anonymous):

12=k/6

OpenStudy (anonymous):

so the answer is b

OpenStudy (mathstudent55):

You are one step ahead of me. All I meant was for you to write \(V = \dfrac{k}{T}\), but you are already at the next step which is to replace V and T with the given values so you can calculate k. \(12 = \dfrac{k}{6} \) Now you need to solve for k.

OpenStudy (mathstudent55):

\(k = 72\) \(V = \dfrac{72}{T}\) You are correct.

OpenStudy (anonymous):

thank you very much.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

and can you help me with some others that i'm going to post?